# Bayesian Hierarchical Probabilistic Forecasting of Intraday Electricity Prices

### v1

@@ 5.2 Evaluation of point estimates
As a first evaluation, we compute the mean absolute error (MAE) of point estimates $\hat{y}$ for all scenarios (a) - (f). As a benchmark model, we use the live IDFull, in line with the proposed weak-form efficiency assumption of the CID market. In Figure 7, we show the MAE of $\hat{y}$ and the live IDFull for all forecast hours of scenarios (a) - (d). These results indicate that with more transaction data available, not only the MAE becomes smaller, also the forecast model tends to beat the live IDFull more often.
@@ 5.3 Spread sign forecasts
Given the probabilistic forecast, more than point estimates can be extracted. A straight forward extraction are the probabilities to observe an IDFull smaller or larger than the DA price, that is the sign of the spread $\mathrm{eod_{d}}(P_{\mathrm{idfull}})-P_{\mathrm{da}}$, which can be estimate as
@@
Also of interest is the rest IDFull, $\mathrm{eod_{d}}(P_{\mathrm{idfull}})-P_{\mathrm{idfull}}$, that is the difference of the end-of-day value of the IDFull to its current (live) value. A positive rest indicates an average increase of electricity prices until gate closure, and a negative sign indicates falling prices. Their probabilities, estimated by
@@
To evaluate these probabilistic sign forecasts, we consider the following estimator for the sign $s$ of the spread, based on a credibility threshold $p_{0}\geq 0.5$,
@@
that is, if we exceed the credibility threshold, we use the forecasted sign implied by $\hat{p}_{\mathrm{ppd}}(y)$, otherwise we use the sign of the live IDFull. For the sign of the rest we use the simpler estimator
@@
Figure 10: The accuracy of the estimators (16) and (17) for the sign of the spread and the rest are shown, together with the IDFull reference for the spread sign, c.f. Figure 9. Here, the scenarios (e) and (f) are used to investigate the difference in performance using OMP and LASSO as feature selection. The accuracy is determined as number of correctly forecasted days across all hours over the total number of forecasts.
@@ 5.4 Probabilistic forecast scores
To directly target reliability, we test the empirical coverage by counting how often the true $\mathrm{eod}_{d}(P_{\mathrm{idfull}})$ was contained in a PI at given credibility $\alpha$. Normalizing and plotting against $\alpha$, we can visually inspect how closely the empirical coverage follows the theoretical expectation given by the unit diagonal. In Figures 11 and 12 we depict the empirical coverage in the left panels for all scenarios. Averaging the deviation from the theoretical diagonal, we obtain the Average Coverage Error (ACE), which is depicted on the right panels of the same Figures 11 and 12. It can be seen that all forecasts tend to be a little overconfident towards higher credibilities, a signature of the highly volatile prices on the German CID market. Using OMP instead of LASSO reduces the ACE by a factor of about 2. An overall comparison between OMP and LASSO in terms of ACE is shown on the left in Figure 13.
@@ 6 Conclusions
We presented a forecast study of the IDFull electricity price index of the German continuous intraday market for the years 2021 and 2022. These years are characterised by unprecedented volatility and have hardly been studied before. We provided details about reproducing price index and statistics published by EPEX Spot using transaction data of the continuous intraday market.
@@
Another aspect discussed in the literature that we take up in our study is the question of feature selection. Comprehensive forecast models that potentially can beat last price benchmarks may draw from a large pool of regressors. These variables typically exhibit strong co-linearities, impeding robust feature selection. In previous works, LASSO has been proven to be successful for feature selection, and even been declared the gold standard. However, LASSO can still exhibit instabilities under strong co-linearities. More robust is Orthogonal Matching Pursuit (OMP), which we suggest as an alternative to LASSO. In our study, we indeed find a clear improvement using OMP instead of using LASSO with a strong statistical significance.
Our model can be developed further in various ways offering plenty possibilities for future research, such as considering different error distributions with long tails and skewness, adding non-linearities to the basic regression approach, and employing ensemble or mixture models.

### v2

@@ 5.2 Evaluation of point estimates
As a first evaluation, we compute the mean absolute error (MAE) of point estimates $\hat{y}$ for all scenarios (a) - (f). As a benchmark model, we use the live IDFull $P_{\mathrm{idfull}}$, in line with the proposed weak-form efficiency assumption of the CID market. In Figure 8, we show the MAE of $\hat{y}$ and the live IDFull for all forecast hours of scenarios (a) - (d). These results indicate that with more transaction data available, not only the MAE becomes smaller, also the forecast model tends to beat the live IDFull more often.
@@ 5.3 Spread sign forecasts
Given the probabilistic forecast, more than point estimates can be extracted. A straight forward extraction are the probabilities to observe an IDFull smaller or larger than the DA price, that is the sign of the spread $\Delta P_{\mathrm{spread}}=\mathrm{eod_{d}}(P_{\mathrm{idfull}})-P_{\mathrm{da}}$, which can be estimated as
@@
Also of interest is the rest $\Delta P_{\mathrm{rest}}=\mathrm{eod_{d}}(P_{\mathrm{idfull}})-P_{\mathrm{idfull}}$, that is the difference of the end-of-day value of the IDFull to its current (live) value. A positive $\Delta P_{\mathrm{rest}}$ indicates an average increase of electricity prices until gate closure, and a negative sign indicates falling prices. The probabilities of these two cases, estimated by
@@
To evaluate these probabilistic sign forecasts, we consider the following estimator for the sign $s$ of $\Delta P_{\mathrm{spread}}$, based on a credibility threshold $p_{0}\geq 0.5$,
@@
that is, if we exceed the credibility threshold, we use the forecasted sign implied by $\hat{p}_{\mathrm{ppd}}(y)$, otherwise we use the sign of the live IDFull. For the sign of $\Delta P_{\mathrm{rest}}$, we use the simpler estimator
@@
Figure 11: The accuracy of the estimators (16) and (17) for the sign of the spread and the rest are shown, together with the IDFull reference for the spread sign, cf. Figure 10. Here, the scenarios (e) and (f) are used to investigate the difference in performance using OMP and LASSO as feature selection. The accuracy is determined as number of correctly forecasted days across all hours over the total number of forecasts.
@@ 5.4 Probabilistic forecast scores
To directly target reliability, we test the empirical coverage by counting how often the true $\mathrm{eod}_{d}(P_{\mathrm{idfull}})$ was contained in a PI at given credibility $\alpha$. Normalizing and plotting against $\alpha$, we can visually inspect how closely the empirical coverage follows the theoretical expectation given by the unit diagonal. In Figures 12 and 13 we depict the empirical coverage in the left panels for all scenarios. Averaging the deviation from the theoretical diagonal, we obtain the Average Coverage Error (ACE), which is depicted on the right panels of the same Figures 12 and 13. It can be seen that all forecasts tend to be a little overconfident towards higher credibilities, a signature of the highly volatile prices on the German CID market. Using OMP instead of LASSO reduces the ACE by a factor of about 2. An overall comparison between OMP and LASSO in terms of ACE is shown on the left in Figure 14.
@@ 6 Conclusions
We presented a forecast study of the IDFull electricity price index of the German continuous intraday market for the years 2021 and 2022. These years are characterised by unprecedented volatility and have hardly been studied before. We provided details about reproducing price indices and statistics published by EPEX Spot using transaction data of the continuous intraday market.
@@
Another aspect discussed in the literature that we address in our study is feature selection. Comprehensive forecast models that potentially outperform last-price benchmarks often draw from a large pool of regressors. These variables typically exhibit strong collinearities, which impede robust feature selection. Previous works have found that LASSO is effective for feature selection and is thus considered the gold standard. However, LASSO can still exhibit instabilities under strong collinearities. A more robust alternative is Orthogonal Matching Pursuit (OMP), which we suggest as an alternative to LASSO. In our study, we find a clear improvement using OMP instead of LASSO, with strong statistical significance.
In summary, the innovative contribution of this work to electricity price forecasting primarily lies in the Bayesian processing of full parameter uncertainty information rather than reducing it to point estimates, along with the probabilistic analysis of the resulting posterior predictive distributions. Additional contributions include the reproduction and live values of all intraday price indices and statistics, and handling the large set of strongly correlated features with OMP and a regularising prior. Lastly, we advocate for probabilistic modelling using TensorFlow Probability.
Our model can be developed further in various ways offering plenty possibilities for future research, such as systematically exploring different error distributions with long tails and skewness, adding non-linearities to the basic regression approach, and employing ensemble or mixture models.